OpenStudy (anonymous):

A salad bar offers 8 choices of toppings for lettuce. In how many ways can you choose 4 or 5 toppings?

OpenStudy (anonymous):

4 toppings = 8*7*6*5= 1680 combinations of toppings 5 toppings = 8*7*6*5*4= 6720 combinations of toppings thats without choosing two of the same topping in one go